# Backdoor-based Explainable AI Benchmark for High Fidelity Evaluation of Attribution Methods

### v1

@@ Abstract
Attribution methods compute importance scores for input features to explain the output predictions of deep models. However, accurate assessment of attribution methods is challenged by the lack of benchmark fidelity for attributing model predictions. Moreover, other confounding factors in attribution estimation, including the setup choices of post-processing techniques and explained model predictions, further compromise the reliability of the evaluation. In this work, we first identify a set of fidelity criteria that reliable benchmarks for attribution methods are expected to fulfill, thereby facilitating a systematic assessment of attribution benchmarks. Next, we introduce a Backdoor-based eXplainable AI benchmark (BackX) that adheres to the desired fidelity criteria. We theoretically establish the superiority of our approach over the existing benchmarks for well-founded attribution evaluation. With extensive analysis, we also identify a setup for a consistent and fair benchmarking of attribution methods across different underlying methodologies. This setup is ultimately employed for a comprehensive comparison of existing methods using our BackX benchmark. Finally, our analysis also provides guidance for defending against backdoor attacks with the help of attribution methods.
Index Terms:
Explainable AI, feature attribution, evaluation benchmark, neural backdoor.
@@ 1 Introduction
Deep learning models have made remarkable strides across diverse domains [1, 2, 3] owing to their capacity to learn intricate representations through numerous parameters. However, the extensive parameter scale that contributes to their success also renders these models less interpretable for their decisions, limiting their applicability in high-stake tasks [4, 5].
In an effort to provide explanations for model predictions, attribution methods [6, 7, 8] ascribe importance scores, referred to as attributions, to the input features. Benefiting from their theoretical foundations [9, 8, 10], attribution methods are extensively employed in explaining model predictions and establishing model trustworthiness [1, 11, 12, 13]. Nevertheless, despite their popularity, the assessment of attribution methods faces challenges given the lack of ground truth. Consequently, the evaluation often relies on alternative techniques such as the removal of input features or the addition of perturbations [14, 15, 16, 17]. However, such methods lead to a phenomenon known as input distribution shift [18], which compromises the fidelity of the evaluation benchmarks. Conversely, efforts have also been made to conjure attribution ground truth by leveraging training annotations [19, 20], generative framework [21] and model Trojaning [22]. However, these methods still encounter challenges in rigorously delineating the attribution ground truth, which impedes their reliability.
@@
In this article, we first provide a set of clear fidelity criteria that eXplainable Artificial Intelligence (XAI) benchmarks should adhere to. We argue that these benchmarks should ensure fidelity to both the explained model and the input, thereby satisfying both functional mapping invariance (§ 3.1-1) and input distribution invariance (§ 3.1-2). Moreover, we stress that attribution benchmarks should offer verifiable ground truth for attributions and sensitive metrics for evaluating them, which we refer to as attribution verifiability (§ 3.1-3) and metric sensitivity (§ 3.1-4), respectively. Our foundational criteria not only set a standard for benchmarking but also facilitate a clear assessment of the existing XAI benchmarks. We then propose a new XAI benchmark (BackX) for attribution methods based on backdoor attacks [23, 24, 25]. Backdoor attacks render a model sensitive to a specific trigger pattern to control their predictions. We propose to leverage this explicit control to systematically establish ground truth attributions for the model. With thorough analysis, we show that the proposed BackX benchmark is capable of offering a more desirable level of assurance for fulfilling the fidelity criteria over the existing benchmarks.
@@
To maintain benchmark fidelity, attribution evaluation must also deal with other confounding factors related to explanation post-processing and the choice of model output used for the explanation. Different choices in this regard eventually yield distinct attributions [26, 27, 28]. Incidentally, these confounders have often led to the development of additional explanatory techniques designed to clarify the explanations themselves [27, 28, 29], which is counter-productive. There is currently a lack of effort in pursuing a consistent framework for attribution evaluation. Through our well-founded BackX benchmark, we reveal distinct properties of different attribution methods using different setup choices, guiding us to a consistent benchmarking setup across different types of attribution methods. To the best of our knowledge, our work is the first to suggest a consistent setup for different types of attribution methods. Using the identified setup, we eventually assess a range of attribution methods using various trigger controls for different manipulated models under diverse Trojaning techniques. A by-product of this comprehensive analysis is unearthing valuable guidance for defense against backdoor attacks using attribution methods. In summary, this article contributes along the following three key aspects.
@@
    It systematically analyzes the attribution benchmark fidelity to identify key criteria for providing reliable foundations to the evaluation of attribution benchmarks.
@@
    It proposes a backdoor-based XAI benchmark (BackX) for attribution methods. The benchmark leverages controllable attributions through model manipulation, fulfilling the identified fidelity criteria. Superior fidelity of this benchmark is established through theoretical analysis.
@@
    It establishes a consistent setup for a transparent assessment of different types of attribution methods, and performs extensive benchmarking of existing methods with BackX using this setup. In the process, it also offers interesting guidance for defending against backdoor attacks with the help of attribution methods.
2 Related Work
2.1 Attribution Methods
To explain model predictions, attribution methods [6, 7] assign importance scores to input features. Deconvnet [7] and Guided Backpropagation [30] utilize deconvolution technique to compute feature importance. CAM [31] introduces global-average pooling during back-propagation, enhancing localization ability without resorting to deconvolution. GradCAM [32] builds on CAM, improving generalization without introducing additional structures. Compared to CAM-based methods, InputGrad [6] calculates gradients with respect to the input for model explanation. SmoothGrad [26] aggregates input gradients across input samples with Gaussian noise, leading to a notable enhancement in localization ability. FullGrad [10] further integrates gradients from model biases, satisfying additional axioms. To guarantee completeness, integrated gradients (IG) [8] integrates gradients from a reference to the input. However, its performance depends on the choice of reference. To improve reference reliability, IG-SG [26], IG-SQ [18] and IG-Uniform [33] employ perturbed inputs as references. EG [13] employs training samples as references to maintain the distribution invariance. Pan et al. [34] calculated class-specific adversarial samples as the reference. Yang et al. [28] recalibrated attributions with valid references. Akhtar and Jalwana [35] improved IG attribution credibility with a better choice of reference and a modified integration process. Wang et al. [27] explained a contrastive output for distinctive attributions, leading to class-contrastive attributions.
2.2 XAI Benchmarks
Numerous model explanation methods have given rise to a multitude of XAI benchmarks. LeRF and MoRF [15] are extended from pixel flipping [14], which perturbs input samples to test output changes. Similarly, insertion and deletion games [16] offer a benchmark for evaluating attributions of a single input sample. However, Hooker et al. [18] revealed that perturbed images cause input distribution shift, making the benchmarking unreliable. Therefore, ROAR [18] and DiffROAR [36] were proposed to retrain models on the perturbed images, ensuring models are learned within the distribution. ROAD [17] and DiffID [37] offer alternative methods to mitigate input distribution shift without necessitating costly model retraining. On the other hand, sensitivity-n [38], SENSMAX and INFD [39] focus on testing the fidelity of attributions by perturbing the input. Adebayo et al. [40] randomized the model parameters and training labels to test the sanity of attributions. Other efforts have also been made to provide ground truth for the estimated attributions. DiFull, DiPart [20] and Pointing Game [19] employ training annotations for attribution evaluation. Khakzar et al. [41] used model-optimized features as proxy for attribution ground truth. Lin et al. [22] benchmarked explaining tools through analyzing the detection of visible triggers in input for Trojaned models. Additionally, Arras et al. [21] employed the controlled VQA framework to generate synthetic images for benchmarking.
2.3 Neural Backdoor
Backdoors alter a model’s predictions by making it sensitive to certain trigger patterns. This framework is relevant because we leverage backdoors in our BackX. Gu et al. [23] introduced BadNet which mislabels and stamps triggers to poison the training samples. Blended attack [24], ISSBA [25] and WaNet [42] are proposed to generate more stealthy sample-specific triggers, achieving invisible poisoning. Adap-Blend [43] further enhances the stealthiness of attack in the latent space. Due to the clear attribution of backdoor triggers in the models trained to achieve it, we leverage neural backdoors to circumvent the lack of ground truth in attribution evaluation. Given the wide applications of attribution methods in backdoor defense [12, 44], we provide guidance for defense against backdoor attacks using attributions as part of this study.
@@ 3 On Benchmark Fidelity
In this section, we first put forth a set of criteria that a reliable explanation benchmark should adhere to. We then compare existing benchmarks on the proposed criteria.
Let us consider an input sample $x\in\mathbb{R}^{n}$ with its label $y\in\mathbb{R}^{c}$ from a dataset $\mathcal{D}$. A classifier denoted as $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{c}$ is parameterized by $\theta$. To explain the model’s prediction $f(x)$, an attribution explaining tool $\phi:\mathbb{R}^{c}\rightarrow\mathbb{R}^{n}$ is used to generate an attribution map $M$, specifically $M=\phi(f(x))$. An XAI benchmark intends to faithfully evaluate the reliability of the explanation $M$. For a quick reference, we also provide a summary of the notations used in the article in Appx. 9.
@@
3.1 Fidelity Criteria
Our fidelity criteria are based on clear observations, hence they are generally intuitively understood. It is worth emphasizing that our intention of providing the criteria is not to establish exhaustive guarantees for benchmark reliability. Due to the abstract nature of the notion of reliability, it is debatable if such a set of guarantees even exists. Here, we aim at identifying clear foundational properties that are intuitively verifiable through observations.
1) Functional Mapping Invariance: Given a model $f$ to be explained, e.g., Fig. 1(a), functional mapping invariance requires that the attribution benchmarking process does not cause a functional mapping shift to the model, see Fig. 1(b). Suppose a perturbation $\zeta$ is imposed on the model $f$ for benchmarking, resulting in $\zeta(f(x))\neq f(x)$. The perturbation $\zeta$ can cause divergent explanations, making it challenging to ascertain whether the misalignment between the explanations and the model’s predictions stems from the unreliable explanatory tools or the functional mapping shift. Thus, functional mapping invariance is crucial for upholding the benchmark fidelity to the explained model.
2) Input Distribution Invariance: Input distribution invariance mandates the constancy of the input distribution $\mathcal{P}_{\mathcal{D}}$ in benchmarking. Assuming $\xi$ represents an input perturbation that leads to input distribution shift (i.e., $\mathcal{P}_{\mathcal{D}}\neq\mathcal{P}_{\xi(\mathcal{D})}$). This shift causes explanation variation from the input samples to the perturbed input samples [18] - see Fig. 1(c). Similar to functional mapping shift, input distribution shift can also result in divergent explanations, obscuring the true cause of misalignment between the explanations and the model’s predictions. Thus, maintaining input distribution invariance is crucial for ensuring the benchmark’s fidelity to the explained samples.
3) Attribution Verifiability: This criterion requires that the correctness of the estimated attributions is verifiable. Given attributions $M=\phi(f(x))$, a faithful benchmark should be able to provide corresponding ground truth attributions $M^{*}$. The ground truth $M^{*}$ is crucial for reliable benchmarking. Fig. 1(d) illustrates the idea where the normal vector from a sample to the decision boundary and its horizontal and vertical components provide the ground truth attributions ($M^{*}=M^{*}_{X}+M^{*}_{Y}$), allowing verification of an estimated attribution ($M=M_{X}+M_{Y}$) in a quantifiable manner.
4) Metric Sensitivity: Metric sensitivity requires that the metric used by the benchmark exhibits sensitivity to the attribution change. Fig. 1(e) shows examples of both sensitive and insensitive metrics. Given the estimated attribution $M$ with its corresponding ground truth $M^{*}$, an insensitive metric, such as the area enclosed by the normal vectors and their components, yields identical results for both $M^{*}$ and $M$. Thereby, failing to accurately quantify the differences between the attributions. In contrast, the difference between the estimated attribution $M^{*}$ and the ground truth $M$ ($\Delta M_{X}+\Delta M_{Y}$) serves as a sensitive metric, demonstrating a faithful sensitivity to the changes in the attributions.
TABLE I: Fidelity criterion fulfillment of XAI benchmarks.
@@ 3.1 Fidelity Criteria
3.2 Fidelity Comparison
In Tab. I, we present a comparative analysis of different benchmarks with regard to their fidelity criteria assurance. Below, we discuss the details of the ratings provided in the table following three categories of the methods based on their underlying techniques.
1) Input Perturbation-based Methods. Since attribution maps are expected to highlight important input features, attribution benchmarks, including MoRF, LeRF [15], Ins.&Del. Games [16], DiffID [37] and ROAD [17], assess attribution methods by iteratively replacing pixels with zero or noise pixels. However, the input change causes an input distribution shift [18], compromising the reliability of evaluation outcomes. To ensure input distribution invariance, ROAR [18] and DiffROAR [36] retrain the model using perturbed input samples to incorporate the perturbed input samples within the training distribution. However, retraining models violate the criterion of functional mapping invariance. Thus, input perturbation-based attribution benchmarks shown in columns 1-3 of Tab. I struggle to ensure invariance to input distribution and the explained model.
2) Sanity and Sensitivity Checks. Sanity and sensitivity checks are originally proposed to test the fidelity of attribution methods for the explained input samples and models. Benchmarks such as sensitivity-n [38], SENSMAX, and INFD [39] aim to assess the sensitivity of attribution methods under input perturbations. While these benchmarks provide valuable assessment targets, these targets are unachievable by attribution methods in practice, resulting only in a partial guarantee of attribution verifiability. Adebayo et al. [40] randomized model parameters and training labels to test the sanity of attributions. However, this sanity check can lead to a functional mapping shift. Similar to the input perturbation-based techniques, sanity and sensitivity checks in columns 4-5 of Tab. I rely on input or model perturbations, making it challenging to simultaneously satisfy both input distribution invariance and functional mapping invariance.
3) Benchmark with Attribution Ground Truth. Other attribution benchmarks have made efforts to establish attribution ground truth for assessing attribution methods. Pointing Game [19], DiFull & DiPart [20] utilize training annotations as attribution ground truth. However, training annotations only offer partial ground truth information for attributions. Khakzar et al. [41] calculated Null Feature without class information through model optimization as the ground truth for attributions. However, there remains a notable absence of evidence substantiating the relationship between the features generated by model optimization and their corresponding attributions. CLEVR-XAI [21] employs a visual question-answering framework to generate synthetic images to provide controlled evaluations of generated objects. However, the presence of shadows from these objects compromises the strict confinement of attribution to the objects’ physical areas. Lin et al. [22] employ neural Trojan for providing attribution ground truth, but their benchmark fails to guarantee that the model prediction change is solely attributable to Trojaned features. In addition to the partial guarantee of attribution verifiability, these attribution benchmarks evaluate the ratio between the defined ground truth region and the region with high attributions, compromising sensitivity in their evaluation metrics for attributions. From Tab. I, it is evident that a full guarantee of attribution verifiability presents a significant challenge.
Overall, our benchmark, discussed in the forthcoming section, stands out as more desirable than the existing techniques on the fidelity criteria.
Figure 2: The pipeline of BackX. Step 1 embeds a backdoor into a benign model by retraining it on a poisoned training set. Step 2 uses the Trojaned model to generate predictions. Step 3 applies attribution methods to explain the predictions. Step 4 recovers a sample from poisoned sample by replacing its pixels from clean sample, as guided by attribution mask. Step 5 assesses attribution methods.
4 BackX Benchmark
In this section, we commence by illustrating the pipeline of our BackX benchmark and then we introduce a set of metrics carefully crafted to evaluate the capabilities of attribution methods. Finally, we provide a discussion on the assurance of benchmark fidelity criteria in the BackX benchmark.
4.1 Benchmark Framework
In Fig. 2, we present an illustration of our BackX benchmark pipeline. Given a clean training set $\mathcal{D}$, we generate a poisoned training set $\mathcal{\tilde{D}}$ by incorporating a trigger $v$ into certain portions of the clean samples and modifying the true label $y$ with the poisoned target label $\tilde{y}$. This poisoned training set is then used to transform a benign model $f$ into a Trojaned model $\tilde{f}$ in Step 1 of Fig. 2. The Trojaned model is employed to generate predictions $\tilde{f}(x)$ and $\tilde{f}(\tilde{x})$ of clean and poisoned samples $x$ and $\tilde{x}$, as shown in Step 2 of Fig. 2. Then, an attribution method $\phi$ explains the output prediction $\tilde{f}_{\tilde{y}}(\tilde{x})$ for class $\tilde{y}$, producing an attribution map $M$ Step 3 of Fig. 2. Subsequently, a mask $S^{(k)}\in\{0,1\}^{n}$ is computed, identifying the top $k$most important input features (e.g., input pixels) as specified by the attribution map. The mask is then utilized to remove the poisoned portion of the sample to construct a recovered sample $\hat{x}$, see Step 4 of Fig. 2. Specifically, a clean patch is extracted from the clean image using the mask. The recovered sample is constructed by replacing the corresponding pixels in the poisoned sample with the extracted patch. Mathematically, the recovered sample $\tilde{x}$ is defined as
$\hat{x}=\tilde{x}\odot(\boldsymbol{1}-S^{(k)})+x\odot S^{(k)}.$ (1)
Here, the recovered sample $\hat{x}$ estimated from a faithful attribution method is expected to be an accurate transformation of the poisoned sample $\tilde{x}$ to its clean counterpart $x$. Thus, attributions $\phi(\tilde{f}(x))$ for the predictions on the clean samples serve as the attribution ground truth for the recovered sample. Driven by the attainability of ground truth, we benchmark attribution methods using the backdoored model, as illustrated in Step 5 of Fig. 2.
@@ 4.1 Benchmark Framework
In BackX, we comprehensively incorporate both visible and invisible trigger patterns for backdoored models through Blend [24] and ISSBA [25] techniques. For the complete control, we enhance the Blend attack with a watermark trigger to ensure poisoned samples align with the original distribution, focusing the backdoored model solely on the trigger region. The watermark trigger is created using the formula
$v^{(\alpha)}=\alpha\cdot v+(1-\alpha)\cdot x\odot S^{*},$ (2)
where $\alpha\in[0,1]$ represents the visibility of the trigger, and $S^{*}\in\{0,1\}^{n}$ is the mask of trigger $v$. ISSBA attack is utilized to generate input-specific invisible triggers. Further details on the backdoored models and triggers used in our experiments are presented in Appx. 11.

### v2

# A Backdoor-based Explainable AI Benchmark for High Fidelity Evaluation of Attributions

@@ Abstract
Attribution methods compute importance scores for input features to explain model predictions. However, assessing the faithfulness of these methods remains challenging due to the absence of attribution ground truth to model predictions. In this work, we first identify a set of fidelity criteria that reliable benchmarks for attribution methods are expected to fulfill, thereby facilitating a systematic assessment of attribution benchmarks. Next, we introduce a Backdoor-based eXplainable AI benchmark (BackX) that adheres to the desired fidelity criteria. We theoretically establish the superiority of our approach over the existing benchmarks for well-founded attribution evaluation. With extensive analysis, we further establish a standardized evaluation setup that mitigates confounding factors such as post-processing techniques and explained predictions, thereby ensuring a fair and consistent benchmarking. This setup is ultimately employed for a comprehensive comparison of existing methods using BackX. Finally, our analysis also offers insights into defending against neural Trojans by utilizing the attributions.
@@ 1 Introduction
Deep learning models have made remarkable strides across diverse domains (He et al., 2016; Goodfellow et al., 2014; Ren et al., 2015) owing to their capacity to learn intricate representations. However, the extensive parameter scale that contributes to their success also renders these models less interpretable for their decisions, limiting their applicability in high-stake tasks (Gade et al., 2019; Tjoa & Guan, 2020; Ali et al., 2023). In an effort to provide explanations for model predictions, attribution methods (Simonyan et al., 2014; Shrikumar et al., 2017; Sundararajan et al., 2017) ascribe importance scores, referred to as attributions, to the input features. Despite their widespread adoption, reliably evaluating the faithfulness of attribution methods remains an open problem due to the lack of high-fidelity benchmarks. Consequently, the evaluation often relies on surrogate strategies such as feature removal and input perturbations (Bach et al., 2015; Samek et al., 2016; Petsiuk et al., 2018; Rong et al., 2022). However, these approaches often induce input distribution shift (Hooker et al., 2019), undermining the fidelity of the evaluation by unintentionally altering the input distribution. Conversely, efforts have also been made to conjure attribution ground truth by leveraging training annotations (Zhang et al., 2018; Rao et al., 2022). However, these benchmarks still encounter challenges in rigorously delineating the attribution ground truth.
@@
In this paper, we first provide a set of clear fidelity criteria that eXplainable Artificial Intelligence (XAI) benchmarks should adhere to. We argue that these benchmarks should ensure fidelity to both the explained model and the input, thereby satisfying both functional mapping invariance and input distribution invariance. Moreover, we stress that attribution benchmarks should offer verifiable ground truth for attributions and sensitive metrics for evaluation, which we refer to as attribution verifiability and metric sensitivity, respectively. Our foundational criteria not only set a standard for benchmarking but also facilitate a clear assessment of the existing XAI benchmarks. We then propose a new XAI benchmark (BackX) for attribution methods based on backdoor neural Trojans (Gu et al., 2019) that render a model sensitive to a specific trigger pattern to control their predictions. We propose to leverage this explicit control to systematically establish ground truth attributions. Through a theoretical analysis, we establish a superior fidelity of our BackX.
@@
To maintain benchmark fidelity, attribution evaluation must also deal with other confounding factors related to post-processing techniques and the explained model’s predictions used for the explanation. Different choices in this regard eventually yield distinct attribution explanations (Smilkov et al., 2017; Wang & Wang, 2022; Yang et al., 2023b). There is currently a lack of effort in pursuing a standardized setup for attribution estimation. Through our well-founded BackX benchmark, we reveal distinct properties of attribution methods using different setup choices, guiding us to a standardized setup across different attribution methods. To the best of our knowledge, our work is the first to suggest a fair setup for existing attribution methods. Using the identified setup, we eventually assess a range of attribution methods using various trigger controls under diverse Trojaning techniques across both vision and language domains. A by-product of this comprehensive analysis is unearthing insights for defense against neural Trojans using attributions. In summary, this article contributes along the following three key aspects.
@@
    It systematically analyzes the attribution benchmark fidelity to identify key criteria for providing reliable foundations for the evaluation of attribution benchmarks.
@@
    It proposes a backdoor-based XAI benchmark (BackX) for attribution methods. BackX leverages controllable attributions through model manipulation, fulfilling the identified fidelity criteria.
@@
    It establishes a standardized setup for a fair assessment of attribution methods. Extensive benchmarking is conducted across both visual and textual domains, uncovering distinct method behaviors and offering guidance for backdoor defense via attribution methods
@@ 2 On Benchmark Fidelity
Let us consider an input sample $x\in\mathbb{R}^{n}$ with its label $y\in\mathbb{R}^{c}$ from a dataset $\mathcal{D}$. A classifier denoted as $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{c}$ is parameterized by $\theta$. To explain the model’s prediction $f(x;\theta)$, an attribution explaining tool $\phi:\mathbb{R}^{c}\rightarrow\mathbb{R}^{n}$ is used to generate an attribution map $M$, specifically $M=\phi(f(x;\theta))$. An XAI benchmark intends to faithfully evaluate the reliability of the explanation $M$. For a quick reference, we also provide a summary of the notations used in App. A.1.
In this section, we first put forth a set of criteria that a reliable explanation benchmark should adhere to. We then compare existing benchmarks on the proposed criteria. While the following sections discuss major related works, a more systematic overview is presented in App. A.2.
2.1 Fidelity Criteria
Functional Mapping Invariance. Given a model $f$ to be explained, e.g., Fig. 1(a), functional mapping invariance requires that the attribution benchmarking process does not cause a functional mapping shift to the model, see Fig. 1(b). Suppose a perturbation $\zeta$ is imposed on the model $f$ for benchmarking, resulting in $f(x;\theta+\zeta)\neq f(x;\theta)$. This can lead to divergent explanations $\phi(f(x;\theta+\zeta))$, depending on $\zeta$; which undermines the benchmarking reliability. Thus, functional mapping invariance is crucial for upholding the benchmark fidelity to the explained model.
Input Distribution Invariance. Input distribution invariance mandates the constancy of the input dataset distribution $\mathcal{P}_{\mathcal{D}}$ in benchmarking. Assuming $\xi$ represents perturbations on the input samples that leads to input distribution shift (i.e., $\mathcal{P}_{\mathcal{D}}\neq\mathcal{P}_{\xi(\mathcal{D})}$) - see Fig. 1(c). This shift causes a variation of explanations for the original distribution $\mathcal{D}$ to the perturbed distribution $\xi(\mathcal{D})$. As a result, input distribution shift can lead to divergent benchmarking results of explanations, obscuring the true cause of misalignment between the explanations and the model’s predictions. Thus, maintaining input distribution invariance is crucial for ensuring the benchmark’s fidelity to the explained samples.
Attribution Verifiability. Attribution verifiability requires that the correctness of estimated attributions can be precisely validated. As illustrated in Fig. 1(d), the ground-truth attributions $M^{*}$ for a given input data point can be derived from the normal vector decomposed along the horizontal and vertical axes, such that $M^{*}=M^{*}_{X}+M^{*}_{Y}$. The availability of the ground truth $M^{*}$ enables precise verification of estimated attributions $M=M_{X}+M_{Y}$. Therefore, the benchmark must provide verifiable ground truth attributions $M^{*}$ to ensure its fidelity to the evaluated attributions.
@@ 2.1 Fidelity Criteria
Metric Sensitivity. Metric sensitivity requires that the metric used by the benchmark exhibits sensitivity to the attribution change. Fig. 1(e) shows examples of both sensitive and insensitive metrics. Given the estimated attribution $M$ with its ground truth $M^{*}$, an insensitive metric, such as the area enclosed by the normal vectors and their components, yields identical results for both $M^{*}$ and $M$. This insensitivity fails to accurately quantify the attribution changes. In contrast, the difference of vector components between the estimated attribution $M$ and the ground truth $M^{*}$, denoted by $\Delta M_{X}$ and $\Delta M_{Y}$, serves as a reliable metric. Therefore, the metric used in the benchmark must be sensitive to attribution changes to ensure its fidelity to the evaluation process.
Table 1: Fidelity criterion fulfillment of XAI benchmarks.
@@
  • $\CIRCLE=\text{Strong (1)}$; $\LEFTcircle=\text{Weak (0.5)}$; $\Circle=\text{No fulfillment (0)}$.
2.2 Fidelity Comparison
In Tab. 1, we present a comparative analysis of different benchmarks with regards to their fidelity criteria assurance. We refer to App. A.4 for the details of the ratings provided in the table. The attribution benchmarks that rely on input perturbations, such as MoRF, LeRF (Samek et al., 2016), and Ins.&Del. Games (Petsiuk et al., 2018), fail to ensure input distribution invariance. ROAR (Hooker et al., 2019) and DiffROAR (Shah et al., 2021) retrain models on perturbed input samples to maintain input distribution invariance. However, they introduce shifts in functional mapping, thereby compromising functional mapping invariance. Efforts by ROAD (Rong et al., 2022) and DiffID (Yang et al., 2023a) to address functional mapping shifts fall short in fully ensuring input distribution invariance. Existing sanity checks (Adebayo et al., 2018; Ancona et al., 2018; Yeh et al., 2019) also test the fidelity of attribution methods under model and input variations. Overall, perturbation-based attribution benchmarks and sanity checks in columns 1-5 of Tab. 1 struggle to simultaneously ensure invariance to both functional mapping and input distribution.
In Tab. 1, it is evident that a full guarantee of attribution verifiability presents a significant challenge. In an effort to establish attribution ground truth for verifiability, Pointing Game (Zhang et al., 2018), DiFull & DiPart (Rao et al., 2022) employ training annotations as attribution ground truth. Lin et al. (2021) use neural Trojan for providing attribution ground truth. CLEVR-XAI (Arras et al., 2022) generates synthetic images to provide controlled evaluations. However, these benchmarks fail to provide fully verifiable attributions. The challenge of offering attribution ground truth also adds complexity to upholding metric sensitivity within such benchmarks. Overall, our benchmark, discussed in the forthcoming section, stands out for its higher fidelity compared to existing techniques.
Figure 2: The pipeline of BackX. Step 1 embeds a backdoor into a benign model by retraining it on a poisoned set. Step 2 uses the Trojaned model to generate predictions. Step 3 explains the predictions via attribution methods. Step 4 recovers a sample from a poisoned sample by replacing its pixels from a clean sample, as guided by the attribution mask. Step 5 assesses attribution methods.
3 BackX Benchmark
In this section, we commence by illustrating the pipeline of our BackX benchmark and then we introduce a set of metrics crafted to evaluate the capabilities of attribution methods. Finally, we provide a discussion on the assurance of benchmark fidelity criteria in the BackX benchmark.
3.1 Benchmark Framework
@@
In BackX, we comprehensively incorporate both visible and invisible trigger patterns for Trojaned models through Blend (Chen et al., 2019) and ISSBA (Li et al., 2021). For the complete control, we enhance the Blend attack with a watermark trigger to ensure poisoned samples align with the original distribution, focusing the Trojaned model solely on the trigger region. The watermark trigger follows the formula $v^{(\alpha)}=\alpha\cdot v+(1-\alpha)\cdot x\odot S^{*}$, where $\alpha\in[0,1]$ is the visibility of the trigger, and $S^{*}\in\{0,1\}^{n}$ is the mask of trigger $v$. ISSBA attack is utilized to generate input-specific invisible triggers. Further details on the Trojaned models used in our experiments are in App. A.7.